You are holding a 35 cm x 42 cm sheet of paper horizontally when a thundercloud produced a vertical electric field of magnitude of 45 kN/C at ground level(a) What is the electric flux through the sheet? (b) What would the flux be if you tilt the sheet of paper by 55°? (c) What would the flux be if you hold the sheet of paper vertically?


1
Expert's answer
2021-03-04T11:50:41-0500

Explanations & Calculations


  • Flux is how many field lines go through a unit area perpendicular to it & it is given in the equation form as

"\qquad\qquad
\begin{aligned}
\small \phi &= \small A\cdot\vec{E}\sin\theta
\end{aligned}" where "\small \theta" is the angle by how much the considered plane is tilted with the electric field.

  • Using this, those needed answers can be found


  • Area of the sheet

"\qquad\qquad
\begin{aligned}
\small A&= \small 0.35m\times0.42m=0.147m^2
\end{aligned}"

a)

  • As the sheet is horizontal, it is perpendicular to the field, then,

"\qquad\qquad
\begin{aligned}
\small \phi_1&= \small 45\times 10^3NC^{-1}\times 0.147m^2\\
&= \small 6615\,Nm^2C^{-1}
\end{aligned}"


b)

  • As the plane is tilted it is "\small (90^0-55^0=)35^0" slanted from the verticle: hence from the field. Then,

"\qquad\qquad
\begin{aligned}
\small \phi_2 &= \small 45\times 10^3\times 0.147m^2\times \sin35\\
&= \small 3794.21\,Nm^2C^{-1}
\end{aligned}"


c)

  • As the plane is kept verticle in this orientation, it makes "\small 0^0" with the field, which means no field line passes through the plane.
  • Hence the flux is zero

"\qquad\qquad
\begin{aligned}
\small \phi_3&= \small 45\times 10^3\times 0.147m^2\times \sin0\\
&= \small 0
\end{aligned}"

  • Always consider the field/resolution of it which is perpendicular to the plane.